OpenStudy (anonymous):

Find the missing coordinate for the given slope and points. m = -1 (1, 2), (2, ) How do you solve this problem? I am stuck. Find the missing coordinate for the given slope and points. m = -1 (1, 2), (2, ) How do you solve this problem? I am stuck. @Mathematics

OpenStudy (karatechopper):

ok so hold on i will draw it

OpenStudy (anonymous):

call the missing one x. so x-2/2-1=-1 x-2/1=-1 multiply by 1 to both sides to get x-2=-1 add 2 to get x=1

OpenStudy (anonymous):

answer is 1

OpenStudy (anonymous):

m = y2 - y1/x2 - x1 replace the given value into equation then you will ge y2 = 1

OpenStudy (karatechopper):